A UK based manufacturer of satellites is reviewing its costs, outputs and profits for the previous year and has approached you to offer some economic analysis. You find the following data: MC = 3Q^2 + 4Q + 2e^2Q+ 7 Assume there are no fixed costs. (a) Determine the total cost function and advise what total costs would be if 10 units are produced. 


1
Expert's answer
2021-01-08T12:57:20-0500

(a) The total cost function is:

"TC = Q^3 + 2Q^2 + e^{2Q} + 7Q + FC,"

where FC is fixed cost.

If 10 units are produced, then the total costs would be:

"TC = 10^3 + 2\u00d710^2 + e^{2\u00d710} + 7\u00d710 + FC = 485,166,465.41 + FC."